OpenStudy (anonymous):

How many combinations of 4 students can a teacher choose from 30 students? A. 657,720 B. 120 C. 810,000 D. 27,405

OpenStudy (anonymous):

you are being asked for "30 choose 4" written as \[\binom{30}{4}\] or sometimes \[_{30}C_4\] or even \[_{30}C_4\] and computed via \[\binom{30}{4}=\frac{\overbrace{30\times 29\times 28\times 27}^{\text {4 terms}}}{4\times 3\times 2}\]

OpenStudy (anonymous):

most calculators have this button so you don't have to do the actual computation, but if you have to do it by hand, since the answer is a whole number, cancel first, multiply last i would do this http://www.wolframalpha.com/input/?i=30+choose+4